# RL in Name Only? Analyzing the Structural Assumptions in RL post-training for LLMs

### v2

@@ 5 Length Bias in GRPO
Building on the length bias proof with GAE estimates in PPO [6], we analyze how GRPO’s group-normalized advantage estimation introduces a similar bias toward longer responses, even though GRPO does not use GAE. For a response with reward $r_{i}$ in a group $\mathcal{G}$, the group relative advantage is defined as:
$A_{i}=\frac{r_{i}-\mu_{\mathcal{G}}}{\sigma_{\mathcal{G}}},$ (10)
where $\mu_{\mathcal{G}}$ and $\sigma_{\mathcal{G}}$ are the mean and standard deviation of rewards in the group.
Correct responses ($r_{i}>\mu_{\mathcal{G}}$, so $A_{i}>0$): For loss calculation, the positive advantage is divided by the response length. Thus, shorter correct responses yield a larger per-token advantage, incentivizing shorter answers.
Incorrect responses ($r_{i}<\mu_{\mathcal{G}}$, so $A_{i}<0$): Similarly, the negative advantage is also divided by the response length. As a result, the longer incorrect responses reduce the penalty applied to each token. This creates an incentive for the model to generate longer incorrect outputs in order to minimize the average per-token penalty. Thus:
  • For correct answers, shorter responses yield higher per-token rewards.
  • For incorrect answers, longer responses reduce the per-token penalty, subtly encouraging longer responses.
As we can see that this behavior results from structural assumptions, particularly, dividing the advantage equally across tokens and scaling it by the length of the response. Yet, the longer responses are often interpreted as signs of ’learning to reason’ [8], even though they result from these assumptions. We will provide further analysis in the Appendix.
@@ 6.1 Experimental Setup
Datasets: We present our analysis on two datasets - 1) the GSM8K dataset [3], which is a widely used benchmark dataset consisting of grade school math problems, designed to evaluate the reasoning capabilities of large language models. It contains 8.5K problems, each paired with a question and a answer. The dataset is divided into 7.5K training problems and 1K test problems. And, 2) Countdown dataset [4], which is a generalized version of the classic 24 Game [17], where the objective is to combine a set of input numbers using basic arithmetic operations (+, –, ×, ÷) to reach a specified target number. In this dataset, each problem consists of 3 to 4 two-digit input numbers, with the target number also being a two-digit number. The dataset contains 9K examples, split into 8K training instances and 1K test instances.
Models: We conduct our experiments using two base models from the Qwen-2.5 family: (1) Qwen-2.5-0.5B and (2) Qwen-2.5-1.5B.
@@
  • Filtered-ISFT: Filtered Iterative Supervised Fine-Tuning using only negative outputs. In this baseline, the base model is trained on filtered negative outputs sampled from the real-time policy by minimizing the log-likelihood of the tokens in these negative outputs.
@@
Training Hyper-parameters: For all baseline implementations, we use the VERL pipeline [13]. All baselines are trained with consistent hyperparameter settings to ensure fair comparison. The training batch size is set to 64 and the mini-batch size to 8 for both datasets. To ensure an equal number of training samples across baselines, we sample 5 responses per question prompt for each method. During response rollouts, the temperature is set to 0.6. The maximum prompt length is set to 512 for GSM8K and 256 for countdown, while the maximum response length is fixed at 1024 for both datasets. The learning rate is set to 1e-6. For GRPO, we set the KL divergence coefficient to $\beta=1e-3$. In the case of Filtered-ISFT+–, we apply a constant weighting of 0.5 to both positive and negative responses throughout training. Model checkpoints are saved every 10 steps for evaluation on the test set. All experiments on the GSM8K dataset are conducted using a single A100 GPU, while Countdown experiments are run using two H100 GPUs. For the GSM8K dataset, both Qwen-2.5-0.5B and Qwen-2.5-1.5B models are trained for 145 global time steps, corresponding to 5 epochs. For the Countdown dataset, Qwen-2.5-0.5B is trained for 600 global time steps and Qwen-2.5-1.5B is trained for 240 global time steps.
@@ 6.2 Results and Discussion
Figure 1 & 2 show the results of baseline comparisons for the Qwen-2.5 models (0.5B and 1.5B parameters) on the GSM8K and Countdown datasets, respectively.
On GSM8K, all baselines show comparable performance. The accuracy of Qwen-2.5-0.5B improves from 0.6% (base model) to 65%, while Qwen-2.5-1.5B improves from 22.7% to 85%. Filtered Filtered-ISFT+–, which incorporates both positive and negative samples, exhibits training dynamics and test accuracy nearly identical to GRPO. Furthermore, using only positive samples, Filtered-ISFT+– has similar performance, suggesting that the core benefits of GRPO can be replicated without reinforcement learning under these structural assumptions.
(a) Base-Model:Qwen-2.5-0.5B
(b) Base-Model:Qwen-2.5-1.5B
Figure 1: Average train and test accuracy on GSM8K dataset
@@
On Countdown dataset, we observe greater variance across baselines. GRPO without KL regularization shows lower performance, possibly due to increased exploration during training. The base model accuracies for Qwen-2.5-0.5B and Qwen-2.5-1.5B are 4.5% and 9.7%, which improve to 45% and 70% respectively after post-training. On validation, Filtered ISFT+ achieves slightly higher performance than GRPO with Qwen-2.5-0.5B, while GRPO yields better results with Qwen-2.5-1.5B. This results also shows that the GRPO and Filtered-ISFT+– has similar performance.
@@
As a sanity check, we also evaluated Filtered-ISFT, which uses only negative samples during fine-tuning. While its performance on GSM8K was on par with other baselines, it significantly underperformed on Countdown. This indicates that negative-only supervision may be inadequate for more challenging tasks like Countdown, where positive examples likely play a critical role in guiding learning and improving generalization.
(a) Base-Model:Qwen-2.5-0.5B
(b) Base-Model:Qwen-2.5-1.5B
Figure 2: Average train and test accuracy on Countdown dataset
@@
Our empirical analysis demonstrates that, under the given structural assumptions, reinforcement learning on current LLM MDP architectures is effectively equivalent to filtered iterative supervised fine-tuning (SFT), incorporating both positive and negative samples.
Figure 3 shows the average response length throughout training on GSM8K and Countdown using Qwen-2.5-0.5B Model and Qwen-2.5-1.5B Model respectively on train-set. We observe that the RL post training with GRPO, initially reduces response length, but it increases again in later stages. We attribute this behavior to RL’s tendency to incentivize longer responses as shown in Section 5, driven by the uniform distribution of advantages across all tokens along with length scaling.
(a) Base-Model:Qwen-2.5-0.5B, GSM8k
(b) Base-Model:Qwen-2.5-1.5B, Countdown
Figure 3: Average response Length over training.
7 Conclusion
@@
In this work, we examine the structural assumptions underlying the popular LLM MDP formulation and analyze their impact on the application of reinforcement learning. Focusing on the widely adopted GRPO algorithm introduced by DeepSeek[8], we show that under assumptions such as representing states as sequences of actions, using terminal rewards, and uniformly distributing credit, GRPO closely resembles filtered iterative supervised fine-tuning. Our empirical results demonstrate that variants of Filtered Iterative SFT with both positive and negative samples, achieve performance comparable to GRPO. We also observe that RL-trained LLMs tend to generate longer responses, an effect often misattributed to improved reasoning. However, our findings suggest that this effect may result from biases in the training setup, specifically due to the uniform distribution of advantages combined with length-based scaling. While our paper shows that the specific structural assumptions used in the LLM-MDP framework popularized by DeepSeek R1 make RL partially degenerate, this doesn’t preclude other LLM-MDP formulations such as the 2 LLM formulation described in [15].
8 Acknowledgements
@@
This research is supported in part by ONR grant N0001423-1-2409, DARPA grant HR00112520016, and gifts from Qualcomm, J.P. Morgan and Amazon. We thank the entire Yochan group for helpful discussions, and Prof. Dimitri Bertsekas and Prof. Angelia Nedic for their insights on reinforcement learning.
@@ A.1 Length bias in Correct Solutions with different think parts
However, as the length of the think part approaches zero, meaning "no think part", the RL objective tends to maximize the conditional probability of solutions without reasoning due to higher advantages. Yet, the initial likelihood of producing correct solutions without think part may be low, particularly for harder or unsolvable problems compared to solvable ones. Consequently, as illustrated in Figure 3 of main paper, we observe that the average response length decreases when the model encounters solvable problems but increases with unsolvable problems.
Thus, the think part emerge based on the initial likelihood and prior probabilities assigned by the base model, influencing the final solution depending on problem difficulty. We demonstrate that reinforcement learning consistently favors shorter correct responses over longer ones. However, we highlight that depending on the difficulty of the problem, it is unlikely the shortest feasible response would entirely lack a reasoning component. Instead, some minimal prompt augmentations or reasoning steps typically persist.
@@ A.3 Length Bias - Experimental Results
Figure 4 presents an empirical analysis of response length and number of correct vs incorrect samples on Countdown dataset during RL post-training of the Qwen-2.5-1.5B model using GRPO. The left side of the figure shows the average response lengths for correct and incorrect samples, while the right side shows the number of correct and incorrect responses throughout training.
These observations support our analysis from Sections A.1 and  A.2. Specifically, the average length of incorrect responses increases as training progresses, whereas the length of correct responses remains relatively stable. Moreover, the number of correct responses consistently rises, while the number of incorrect responses declined. Thus, we analyze the increase in average response length by separating it into contributions from correct and incorrect responses. Our findings show that the increase is mainly caused by the incorrect responses. This can be explained by the structural assumptions, particularly by uniform advantage distribution across tokens and scaling the advantage by length during training. Moreover, we find no evidence that longer responses lead to better reasoning and that the reasoning abilities could emerge due to increased response length.
Appendix B Additional Experiments Results
Figures 5 and 6 illustrate the average train and test accuracies of all baselines across 2 random seeds [^2] Note: Due to the high computational cost, all experiments in this analysis are conducted using only two random seeds for statistical evaluation. for both Qwen-2.5 base models evaluated on the GSM8K and Countdown datasets, respectively.
(a) Base-Model:Qwen-2.5-0.5B
(b) Base-Model:Qwen-2.5-1.5B
Figure 5: Performance comparison of all baselines (except Filtered-ISFT, omitted due to high variance) on GSM8K train and test datasets for Qwen-2.5-0.5B (left column) and Qwen-2.5-1.5B (right column). The y-axis represents the average test and train accuracy computed over 2 random seeds and the shaded region indicate the $95\%$ confidence interval.
For GSM8K, all baselines exhibit comparable performance on both the train and test datasets. For the Countdown dataset, with the Qwen-2.5-0.5B model, GRPO-wo-KL shows slightly higher performance compared to other methods, while GRPO and Filtered-ISFT+– achieve similar results. With the Qwen-2.5-1.5B model, GRPO demonstrates higher performance; however, when considering the $95\%$ confidence intervals, GRPO and Filtered-ISFT+– effectively perform equivalently.
(a) Base-Model:Qwen-2.5-0.5B
(b) Base-Model:Qwen-2.5-1.5B
Figure 6: Performance comparison of all baselines (except Filtered-ISFT, omitted due to high variance) on Countdown train and test datasets for Qwen-2.5-0.5B (left column) and Qwen-2.5-1.5B (right column). The y-axis represents the average test and train accuracy computed over 2 random seeds and the shaded region indicate the $95\%$ confidence interval.

### v3

@@ 5 Length Bias in GRPO
We observe that applying GRPO within the LLM-MDP framework induces a length bias in the learned model. Our empirical analysis further demonstrates that training on incorrect responses significantly contributes to this effect. Given a response with reward $r_{i}$ in a group $\mathcal{G}$, the group relative advantage is defined as: $A_{i}=\frac{r_{i}-\mu_{\mathcal{G}}}{\sigma_{\mathcal{G}}},$ where $\mu_{\mathcal{G}}$ and $\sigma_{\mathcal{G}}$ are the mean and standard deviation of rewards in the group.
Correct responses ($r_{i}>\mu_{\mathcal{G}}$, so $A_{i}>0$): Distributing the positive advantage equally across tokens results in higher per-token rewards for shorter outputs, thereby incentivizing brevity.
Incorrect responses ($r_{i}<\mu_{\mathcal{G}}$, so $A_{i}<0$): Distributing the negative advantage over more tokens lowers the per-token penalty, which encourages the generation of longer incorrect outputs.
This length bias arises from structural assumptions in the learning framework, specifically, dividing the advantage uniformly across tokens and scaling it with response length. We empirically show that this bias also appears in F-ISFT+-, similar to GRPO. Notably, F-ISFT+-, which is trained on both correct and incorrect responses, produces longer outputs than F-ISFT+, which is trained only on positive samples. Yet, the longer responses are often interpreted as signs of ’learning to reason’ (guo2025deepseek), even though they arise from disproportionately longer incorrect responses caused by structural assumptions in the degenerate LLM-MDP framework. We also provide further analysis on length bias in the Appendix A.
@@ 6.1 Experimental Setup
Datasets: We present our analysis on two datasets - 1) the GSM8K dataset (cobbe2021gsm8k), which is a widely used benchmark dataset consisting of grade school math problems, designed to evaluate the reasoning capabilities of large language models. It contains 8.5K problems, each paired with a question and an answer. The dataset is divided into 7.5K training problems and 1K test problems. And, 2) Countdown dataset (wikipediaCountdowngame), which is a generalized version of the classic 24 Game (yang2022chain), where the objective is to combine a set of input numbers using basic arithmetic operations (+, –, ×, ÷) to reach a specified target number. In this dataset, each problem consists of 3 to 4 two-digit input numbers, with the target number also being a two-digit number. The dataset contains 9K examples, split into 8K training instances and 1K test instances. We also report the accuracy on the MATH-500 dataset, a 500-problem test set evaluated using models trained on GSM8K.
Models: We conduct our experiments using two base models from the Qwen-2.5 family: (1) Qwen-2.5-0.5B and (2) Qwen-2.5-1.5B. and two base models from the Llama-3.2 family: (1)Llama-3.2-1B-Instruct and (2) Llama-3.2-3B-Instruct.
@@
  • Filtered-ISFT-: Filtered Iterative Supervised Fine-Tuning using only negative outputs. In this baseline, the base model is trained on filtered negative outputs sampled from the real-time policy by minimizing the log-likelihood of the tokens in these negative outputs.
@@
Training Hyper-parameters: For all baseline implementations, we use the VERL pipeline (sheng2024hybridflow). All baselines are trained with consistent hyperparameter settings to ensure fair comparison. The training batch size is set to 64 and the mini-batch size to 8 for both datasets. To ensure an equal number of training samples across baselines, we sample 5 responses per question prompt for each method. During response rollouts, the temperature is set to 0.6. The maximum prompt length is set to 512 for GSM8K and 256 for countdown, while the maximum response length is fixed at 1024 for both datasets. The learning rate is set to 1e61e-6. For GRPO, we set the KL divergence coefficient to $\beta=1e-3$. In the case of Filtered-ISFT+–, we apply a constant weighting of 0.5 to both positive and negative responses throughout training. Model checkpoints are saved every 10 steps for evaluation on the test set. All experiments on the GSM8K dataset are conducted using a single A100 GPU, while Countdown experiments are run using two H100 GPUs. For the GSM8K dataset, both models of the Qwen and Llama family are trained for $145$ global time steps, corresponding to $5$ epochs. For the Countdown dataset, Qwen-2.5-0.5B is trained for 600 global time steps and Qwen-2.5-1.5B &\& Llama-3.2-3B-Instruct are trained for $300$ global time steps.
@@ 6.2 Results and Discussion
Performance of GRPO vs. F-ISFT variants – Table 1 shows the performance comparisons of all baselines for both Qwen-2.5 models (0.5B and 1.5B) on the GSM8K and Countdown datasets. On GSM8K, all baselines achieve comparable performance, and both Qwen-2.5-0.5B and Qwen-2.5-1.5B show substantial gains after post-training. Filtered-ISFT+-, which incorporates both positive and negative examples, exhibits training dynamics and test performance closely aligned with GRPO. Similarly, Filtered-ISFT+, trained only on positive samples, achieves performance comparable to both GRPO and Filtered-ISFT+-. On the Countdown dataset, greater variance is observed across baselines. GRPO without KL regularization shows lower performance, possibly due to increased exploration during training. While both base models start with low accuracy, post-training significantly improves their performance. On the smaller model, Filtered-ISFT+ slightly outperforms GRPO, whereas on the larger model, GRPO achieves slightly better results.
Methods Qwen-2.5-0.5B Qwen-2.5-1.5B
GSM8k CD8k GSM8k CD8k
Base Model $00.07$ $00.05$ $22.67$ $00.09$
GRPO $\mathbf{55.87}$ $37.73$ $\mathbf{78.24}$ $\mathbf{71.43}$
GRPO-wo-KL $55.19$ $42.43$ $76.87$ $70.82$
Filtered-ISFT+ $51.71$ $\mathbf{44.01}$ $74.98$ $53.57$
Filtered-ISFT- $55.27$ $00.00$ $76.04$ $00.00$
Filtered-ISFT+- $55.72$ $37.86$ $76.37$ $65.89$
Table 1: Performance of all baselines on test dataset of GSM8k and CD8k for both Qwen-2.5 family models (0.5B and 1.5B). For each dataset, the highest score is highlighted in bold.
@@
As a sanity check, we also evaluated Filtered-ISFT–, which uses only negative samples during fine-tuning. While its performance on GSM8K was on par with other baselines, it significantly underperformed on Countdown. This suggests that negative-only supervision may be inadequate for more challenging tasks such as Countdown, where positive examples likely play a critical role in guiding learning and improving generalization.
Methods Llama-3.2-1B-Instruct Llama-3.2-3B-Instruct
GSM8k Math-500 CD8k GSM8k Math-500 CD8k
Base Model $35.78$ $23.40$ $0.00$ $75.13$ $40.80$ $00.01$
GRPO $62.01$ $\mathbf{33.02}$ $0.00$ $\mathbf{84.59}$ $34.00$ $55.96$
Filtered-ISFT+ $57.31$ $25.40$ $0.00$ $81.95$ $48.60$ $\mathbf{56.92}$
Filtered-ISFT+- $\mathbf{63.07}$ $30.20$ $0.00$ $83.54$ $\mathbf{49.40}$ $54.91$
Table 2: Performance of all baselines on the test sets of GSM8k, MATH-500, and CD8k for the LLaMA-3.2 Instruct models (1B and 3B). For each dataset, the highest score is highlighted in bold. Note that results on the MATH-500 dataset are obtained using models trained on the GSM8k dataset.
We further evaluated larger models from the Llama-3.2 Instruct family (1B and 3B) on GSM8K, Countdown, and MATH500, with results reported in Table 2. Experiments with Llama-3.2-1B-Instruct and Llama-3.2-3B-Instruct show that F-ISFT+- and GRPO perform similarly on GSM8K and CD8K. On MATH500, GRPO performs better on the 1B model, while F-ISFT+- achieves slightly higher performance on the 3B model. Although F-ISFT+ has comparatively lower performance on the 1B model, all baselines achieve similar performance across datasets on the 3B model.
@@
These results demonstrate that GRPO and Filtered-ISFT+- perform similarly. Thus, our empirical analysis show that under the given structural assumptions, RL in the current LLM-MDP framework is effectively equivalent to F-ISFT incorporating both positive and negative samples.
Length bias analysis - The figures in 1 present results on the Countdown test dataset at different evaluation steps during the post-training of the Qwen-2.5-1.5B model. The left panel shows the number of correct and incorrect responses, where we observe that the number of correct responses increases steadily while the number of incorrect responses decreases across all baselines. The right panel shows the average token response length for both correct and incorrect responses. For correct responses, we find a slight increase in the average response length for GRPO and Filtered-ISFT+-, while it remains relatively constant for Filtered-ISFT+. For incorrect responses, the average length decreases initially but then increases substantially for GRPO and Filtered-ISFT+-, whereas Filtered-ISFT+ maintains comparable response lengths for both correct and incorrect responses.
Figure 1: Left: Number of correct and incorrect responses. Right: Average response lengths for correct and incorrect responses, on the Countdown test dataset at each evaluation step during post-training of the Qwen-2.5-1.5B base model using GRPO and variant of Filtered-ISFT algorithm.
These length-based analyses indicate that Filtered-ISFT+- also incentivizes longer outputs, often producing even longer responses than Filtered-ISFT+. This suggests that the inclusion of incorrect responses during training contributes to increased output length, similar to the behavior observed in GRPO. We attribute this to RL’s inherent bias toward longer responses as discussed in Section 5 which arises from the uniform distribution of advantages across tokens combined with length scaling. Empirically, we observe that training on incorrect responses leads to increased response length, as the loss continues to decrease with increasing response length on these examples.
7 Discussion and Conclusion
@@
In this work, we critically examined the structural assumptions underlying the popular LLM-MDP formulation and analyzed how they undermine RL applications to LLM post-training. Focusing on the widely adopted GRPO algorithm introduced by DeepSeek (guo2025deepseek), we showed that by representing states as nothing more than sequences of actions, applying only terminal rewards, and uniformly distributing credit, GRPO closely resembles filtered iterative supervised fine-tuning. Our empirical results further demonstrated that variants of filtered iterative SFT, leveraging both positive and negative samples, achieve performance comparable to GRPO. We also observed that RL-trained LLMs tend to generate longer responses, an effect often misattributed to improved reasoning capabilities. Our findings suggest that this phenomenon actually arises from biases in the training setup, specifically the uniform distribution of advantages across tokens combined with length-based scaling effects.
A key implication of our analysis is that misunderstanding the consequences of these structural assumptions has led to a number of downstream research directions that attempt to correct for symptoms of the formulation, rather than addressing the root cause. For example, recent works have proposed explicitly optimizing for shorter reasoning traces by introducing length penalties in the reward function (arora2025training; team2025kimi) or by avoiding training on longer incorrect responses indirectly, e.g. by sorting the responses by length and correctness  (shrivastava2025sample). While these methods appear to mitigate the tendency of RL-trained LLMs to produce excessively long outputs, they do so without recognizing that such length biases are intrinsic to the degenerate reward allocation mechanism of the current LLM-MDP framework. As a result, these interventions risk obscuring the fundamental issue and may unintentionally hinder the development of more principled approaches.
A promising direction may be to revisit the MDP formulation itself. The 2-LLM framework described in (valmeekamsystematic) provides one such avenue. In this formulation, a chain-of-thought (CoT) LLM policy generates intermediate reasoning steps as actions conditioned on the problem prompt, while a main LLM generates the final solution conditioned on both the prompt and the CoT actions. This setup more faithfully reflects the sequential decision-making structure assumed in RL and enables the design of reward functions that meaningfully shape intermediate reasoning. By adopting such alternatives, the full expressive capacity of RL can be more effectively leveraged in guiding LLMs toward better reasoning strategies.
To conclude, our critique is directed at the degenerate LLM-MDP formulation rather than at RL itself. Under the current restrictive assumptions, the potential of RL cannot be fully realized. We hope that our analysis motivates exploration of alternative formulations that can better unlock the benefits of RL for LLM post-training.
Acknowledgment
@@
This research is supported in part by grants from ONR (N00014-25-1-2301 and N00014-23-1-2409), DARPA (HR00112520016), DoD RAI (via CMU subcontract 25-00306-SUB-000), an Amazon Research Award, and a generous gift from Qualcomm. We thank Professors Dimitri Bertsekas and Angelia Nedic for their helpful discussions.
@@ A.1 Length bias in Correct Solutions with different think parts
However, as the length of the think part approaches zero, meaning "no think part", the RL objective tends to maximize the conditional probability of solutions without reasoning due to higher advantages. Yet, the initial likelihood of producing correct solutions without think part may be low, particularly for harder or unsolvable problems compared to solvable ones. Consequently, as illustrated in Figure 2, we observe that the average response length decreases when the model encounters solvable problems but increases with unsolvable problems.
Thus, the think part emerge based on the initial likelihood and prior probabilities assigned by the base model, influencing the final solution depending on problem difficulty. We demonstrate that RL consistently favors shorter correct responses over longer ones. However, we highlight that depending on the difficulty of the problem, it is unlikely the shortest feasible response would entirely lack a reasoning component. Instead, some minimal prompt augmentations or reasoning steps typically persist.
@@ A.3 Length Bias - Experimental Results
Figure 2 presents an empirical analysis of response length and number of correct vs incorrect samples during RL post-training of the Qwen-2.5-1.5B model using GRPO. The left side of the figure shows the average response lengths for correct and incorrect samples, while the right side shows the number of correct and incorrect responses throughout training.
These observations support our analysis from Sections A.1 and  A.2. Specifically, the average length of incorrect responses increases as training progresses, whereas the length of correct responses remains relatively stable. Moreover, the number of correct responses consistently rises, while the number of incorrect responses declined. Thus, we analyze the increase in average response length by separating it into contributions from correct and incorrect responses. Our findings show that the increase is mainly caused by the incorrect responses. This can be explained by the structural assumptions, particularly by uniform advantage distribution across tokens and scaling the advantage by length during training. Moreover, we find no evidence that longer responses lead to better reasoning and that the reasoning abilities could emerge due to increased response length